OpenStudy (anonymous):

918 is 170% of what number show work and do it the "normal" way

OpenStudy (amistre64):

1.70N = 918

OpenStudy (amistre64):

N = 918/1.7

OpenStudy (anonymous):

if you put 170 % in decimal form its 1.7. x*1.7 = 918 x = 918/1.7

OpenStudy (amistre64):

what is a "normal" way ?

OpenStudy (amistre64):

N/100 = 1% N/100 * 170 = 170% N/100 * 170 = 918

OpenStudy (unklerhaukus):

* N/100=N% N/100 * 170 = 170% of N

OpenStudy (unklerhaukus):

i am a physicist so dimension is important